OpenStudy (anonymous):

Once again, get medal :3 Solve for x. 2 • 3 – 2 • 11 = 2(x – 11) A. 2 B. 6 C. 3 D. 9 I know one thing for sure. That 2 • 3 – 2 • 11 is -16, just got to figure out the next part for me :)

OpenStudy (lieutenantgeneral):

x = 3

OpenStudy (anonymous):

6 - 22 = 2x - 22 2x = 6 x = 3

OpenStudy (texaschic101):

2 * 3 - 2 * 11 = 2(x - 11) -- parenthesis first 2 * 3 - 2 * 11 = 2x - 22 -- multiplication next 6 - 22 = 2x - 22 -- add 22 to each side 6 = 2x -- divide by 2 6/2 = x 3 = x

OpenStudy (anonymous):

@texaschic101 @Link360 Thanks for explaining this problem to me, much easier for me to understand *hug*

OpenStudy (texaschic101):

anytime....order of operations can sometimes get confusing :)

OpenStudy (anonymous):

no problem